Find the solution to the equation log of x base 6+ log of (x+5) base 6=2
Check for extraneous solutions

Answer by edjones(7311)

log[6]x+log[6](x+5)=2
6^(log[6]x) * 6^(log[6](x+5))=6^2
x(x+5)=36
x^2+5x-36=0
(x+9)(x-4)=0
x=4 This is the only answer because there can be no log of a neg number.
.
Ed